Katty needs to purchase milk to make ice cream. The supplier charges a $60 fixed fee per order and $4.75 per kg. Her annual holding cost is 25%. She needs 1000 kgs of milk per day. Assume 260 working days per year. If Katty wants to minimize inventory holding and ordering costs, what is the average flow time?

**I know that the correct answer is 2.5 days but how do I get this?
Answer #1

Average flow rate per day(R) = 1000 kgs

Number of days per year = 260 days

Annual demand(D) = R × number of days per year = 1000 × 260 = 260000 kgs

Ordering cost (S) = $60

Holding cost(H) = 25% of cost = 25% of $4.75 = $1.18

Economic order quantity = √(2DS/H)

= √[(2 × 260000 × 60) /1.18]

= √(31200000/1.18)

= 5142.05 KGS

Average inventory = Economic order quantity/2 = 5142.05/2 = 2571.025

Average flow time = Average inventory/average flow rate

= 2571.025/1000

= 2.57 days